OpenStudy (anonymous):

Solve by the substitution method 15x-20y=-1 15y=1+10x

OpenStudy (anonymous):

Minus the 10x to the other side 15x-20y=-1 -10x+15y=1

OpenStudy (anonymous):

Then you will have to get rid of one of the variables so get rid of the x then 2(15x-20y=-1) 3(-10x+15y=1 So it will be 30x-40y=-2 -30x+45y=3 = 5y=1

OpenStudy (anonymous):

Then you will end up as a fraction y=1/5 Plug it into one of the equations -10x+15(1/5)=1

OpenStudy (anonymous):

Then you will end up with -10x=-2 x=2/10 which can be simplified x=1/5 & y=1/5

OpenStudy (anonymous):

Let me check the answers...

OpenStudy (anonymous):

Yes it is correct

OpenStudy (anonymous):

Hello...

OpenStudy (anonymous):

Thanx